# Understanding Users’ Security and Privacy Concerns and Attitudes Towards Conversational AI Platforms

### v1

@@ 1 Introduction
Large language models (LLMs) [1, 2] have revolutionized artificial intelligence (AI), driving the widespread adoption of conversational AI platforms (e.g., ChatGPT [3], Gemini [4], and Claude [5]). These platforms allow users to interact with LLM-based AI systems through different modalities, such as text and voice. They are being integrated into a diverse range of applications, including customer service [6, 7, 8], personal assistants [9, 10], healthcare [11, 12], and finance [13, 14, 15], due to their advanced natural language abilities, enabling more intuitive user interactions.
@@ 2 Related Work
Users’ S&P Concerns and Attitudes. Prior research has explored users’ security and privacy (S&P) concerns [35], behaviors [36], preferences [37, 38], and attitudes [39]. Studies have also explored users’ S&P attitudes and concerns toward specific platforms and technologies. For example, studies on smart homes and IoT devices have highlighted user apprehensions about data collection and device security [40, 41, 42, 43, 44, 45]. Similarly, studies on web browsers and online social networks have investigated users’ views and satisfaction with privacy settings, uncovering a range of concerns about data sharing and control [46]. While prior research has examined general S&P attitudes or concerns toward specific technologies, we focus on the unique challenges and concerns that arise from the human-like interactions facilitated by conversational AI.
@@
While these studies provide valuable initial insights, they have three main limitations. First, they primarily involve participants from the United States, which may not represent the global user base of conversational AI platforms. Privacy concerns and regulatory landscapes vary across jurisdictions (e.g., the GDPR [62] in Europe), leading to differing user concerns and behaviors. Second, these works focus on end-user interactions, overlooking S&P concerns from developers and enterprises deploying these platforms. Lastly, they explore users’ perspectives in a controlled context at a single point in time, failing to capture how user concerns and attitudes evolve in response to new features and emerging risks.
@@ 3.1 Identifying S&P-related posts
To conduct our analysis, we gathered all posts from r/ChatGPT from its inception in December $2022$ till July $2024$, using publicly available Reddit dumps [64]. After removing deleted posts, bot-generated content  [^1] Bot-generated posts are identified using known bot signatures such as the presence of the word “bot” in the username or phrases like “Beep boop, I’m a bot” [65]., and duplicate entries, our final dataset included $\sim 2.5$M posts ($\sim 180$K submissions and $\sim 2.35$M comments, collectively termed “posts”).
@@
To ensure balanced representation and reduce bias towards frequently mentioned keywords, we grouped our keywords into $6$ thematic categories based on prior works [68, 72, 73]. We then performed stratified random sampling, selecting $100$ posts from each group, resulting in $600$ posts potentially related to S&P. To enable the classifier to distinguish S&P-related content from other topics and to capture themes outside our keyword list, we complemented this sample with $600$ randomly selected posts that did not contain any keywords. In total, our seed corpus consisted of $1,200$ posts, balanced between submissions and comments.
@@ 4 RQ1: Users’ S&P Concerns
Our analysis reveals a range of recurring security and privacy (S&P) concerns regarding conversational AI platforms (e.g., ChatGPT). We categorize these concerns into six main themes: ($1$) data collection, ($2$) data usage, ($3$) data retention, ($4$) security vulnerabilities, ($5$) legal compliance, and ($6$) transparency and control. Figure 2 presents the prevalence of each theme and its sub-themes. These prevalence metrics can guide platforms in prioritizing improvements by highlighting the most common user concerns e.g., the need for greater transparency and control around data collection and model training. Below, we detail our findings for each of these themes.
@@ 4.4.2 AI-enabled Applications
Custom GPTs. The recent introduction of customizable GPTs [88], which users can configure with custom instructions, data, and API access and make available to other users via the GPT store [89], has also fueled concerns among users. Users are concerned that adversaries could manipulate these custom GPTs to expose private data (e.g., custom instructions). For example, one user noted: “Even after telling my custom GPT to not leak the information from the knowledge database and the custom instructions through which it is trained, it’s still revealing those things.” Another user described how custom GPTs can be tricked into “spilling secrets” including “sensitive documents”.
@@ 4.4.3 Security of AI-generated code
Although several users acknowledge the utility provided by many conversational AI platforms (e.g., ChatGPT) that enable code generation, debugging, and even direct code execution through built-in interpreters, we find that many developers and professionals express concerns about security vulnerabilities present in the generated code, describing it as “bug filled”. They worry that the code generated by these platforms does not “follow up to date security principles” and many recommend against its use in software systems that handle sensitive data. For example, one user shared: “[I am scared] at the idea of some startup using [AI-generated code] excessively for something that stores and handles sensitive data.” Similarly, a security professional expressed his concern about the scale at which AI-generated code is deployed, stating that “Society is absolutely unprepared for the knock-on effects of releasing all this [insecure] code into production … there are just not enough of us [security researchers] to handle this.”
@@ 7.2.2 For Platforms
Educate Users.
Prior research in S&P has shown that even motivated users often struggle to make safe choices without clear, accessible guidance [112]. Building on these insights, we recommend that platforms develop educational resources to help users understand their operations and data-handling practices. These resources (e.g., tutorials, FAQ pages) can address common misconceptions, provide guidance on best practices for protecting personal data (e.g., adjusting privacy settings), and explain how user data is used for model training. By providing clear, targeted education, platforms can foster a more informed user base.
@@ 7.3 Limitations and Future Work
Given our findings reflect specific concerns from professionals, developers, and other users using these platforms for specific applications, future work could explore users’ S&P concerns and attitudes in specific contexts and domains. By examining sector-specific issues, we can identify unique challenges and develop targeted recommendations for users and stakeholders in those fields. Similarly, future research should investigate the influence of geopolitical and cultural factors on user attitudes toward conversational AI platforms to understand regional differences and develop culturally sensitive solutions and policies.
@@ Appendix B Keywords for Seed Corpus Creation
Table III shows our complete list of keywords for filtering posts. We applied regex-based matching to capture phrase variations (e.g., r’delete(?:my|your)?data’ to match “delete my data” and “delete your data”). Certain terms like access were matched as standalone words to avoid unrelated results (e.g., accessibility).
@@ Appendix C S&P Classification Results
Appendix D S&P-Related Posts Over Time
Figure 4 shows the weekly volume of S&P-related posts over time. On average, there are $344$ S&P-related posts per week.
Figure 4: Weekly counts of S&P-related posts.
TABLE III: Keywords used to filter candidate posts for manual annotation.
@@ Appendix D S&P-Related Posts Over Time
Figure 5: Key events in the development and adoption of conversational AI platforms, including product and feature releases (green), policy and regulatory changes (yellow), corporate (navy), and privacy and security events (red).
@@ Appendix E Multi-Class Classification Results
Table IV shows the performance of the multi-class classification using GPT-4o [75].

### v2

@@ 1 Introduction
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. Published in the IEEE Symposium on Security and Privacy (S&P), 2025. DOI: https://doi.ieeecomputersociety.org/10.1109/SP61157.2025.00241
Large language models (LLMs) [1, 2] have revolutionized artificial intelligence (AI), driving the widespread adoption of conversational AI platforms (e.g., ChatGPT [3], Gemini [4], and Claude [5]). These platforms allow users to interact with LLM-based AI systems through different modalities, such as text and voice. They are being integrated into a diverse range of applications, including customer service [6, 7, 8], personal assistants [9, 10], healthcare [11, 12], and finance [13, 14], due to their advanced natural language abilities, enabling more intuitive user interactions.
@@ 2 Related Work
Users’ S&P Concerns and Attitudes. Prior research has explored users’ security and privacy (S&P) concerns [34], behaviors [35], preferences [36, 37], and attitudes [38]. Studies have also explored users’ S&P attitudes and concerns toward specific platforms and technologies. For example, studies on smart homes and IoT devices have highlighted user apprehensions about data collection and device security [39, 40, 41, 42, 43]. Similarly, studies on web browsers and online social networks have investigated users’ views and satisfaction with privacy settings, uncovering a range of concerns about data sharing and control [44]. While prior research has examined general S&P attitudes or concerns toward specific technologies, we focus on the unique challenges and concerns that arise from the human-like interactions facilitated by conversational AI.
@@
While these studies provide valuable initial insights, they have three main limitations. First, they primarily involve participants from the United States, which may not represent the global user base of conversational AI platforms. Privacy concerns and regulatory landscapes vary across jurisdictions (e.g., GDPR [60] in Europe), leading to differing user concerns and behaviors. Second, these works focus on end-user interactions, overlooking S&P concerns from developers and enterprises deploying these platforms. Lastly, they explore users’ perspectives in a controlled context at a single point in time, failing to capture how user concerns and attitudes evolve in response to new features and emerging risks.
@@ 3.1 Identifying S&P-related posts
To conduct our analysis, we gathered all posts from r/ChatGPT from its inception in December $2022$ till July $2024$, using publicly available Reddit dumps [62]. After removing deleted posts, bot-generated content  [^1] Bot-generated posts are identified using known bot signatures, such as the presence of the word “bot” in the username or phrases like “Beep boop, I’m a bot” [63]., and duplicate entries, our final dataset included $\sim 2.5$M posts ($\sim 180$K submissions and $\sim 2.35$M comments, collectively termed “posts”).
@@
To ensure balanced representation and reduce bias toward frequently mentioned keywords, we grouped our keywords into $6$ thematic categories based on prior works [66, 70, 71]. We then performed stratified random sampling, selecting $100$ posts from each group, resulting in $600$ posts potentially related to S&P. To enable the classifier to distinguish S&P-related content from other topics and to capture themes outside our keyword list, we complemented this sample with $600$ randomly selected posts that did not contain any keywords. In total, our seed corpus consisted of $1,200$ posts, balanced between submissions and comments.
@@ 4 RQ1: Users’ S&P Concerns
Our analysis reveals a range of recurring security and privacy (S&P) concerns regarding conversational AI platforms (e.g., ChatGPT). We categorize these concerns into six main themes: ($1$) data collection, ($2$) data usage, ($3$) data retention, ($4$) security vulnerabilities, ($5$) legal compliance, and ($6$) transparency and control. Figure 2 presents the prevalence of each theme and its sub-themes. These prevalence metrics can guide platforms in prioritizing improvements by highlighting the most common user concerns e.g., the need for greater transparency and control around data collection and model training. Below, we detail our findings for each theme.
@@ 4.4.2 AI-enabled Applications
Custom GPTs. The recent introduction of customizable GPTs [86], which users can configure with custom instructions, data, and API access and make available to other users via the GPT store [87], has also fueled concerns among users. Users are concerned that adversaries could manipulate these custom GPTs to expose private data (e.g., custom instructions). For example, one user noted: “Even after telling my custom GPT to not leak the information from the knowledge database and the custom instructions through which it is trained, it’s still revealing those things.” Another user described how custom GPTs can be tricked into “spilling secrets”, including “sensitive documents”.
@@ 4.4.3 Security of AI-generated code
Although several users acknowledge the utility provided by many conversational AI platforms (e.g., ChatGPT) that enable code generation, debugging, and even direct code execution through built-in interpreters, we find that many developers and professionals express concerns about security vulnerabilities present in the generated code, describing it as “bug filled”. They worry that the code generated by these platforms does not “follow up to date security principles”, and many recommend against its use in software systems that handle sensitive data. For example, one user shared: “[I am scared] at the idea of some startup using [AI-generated code] excessively for something that stores and handles sensitive data.” Similarly, a security professional expressed his concern about the scale at which AI-generated code is deployed, stating that “Society is absolutely unprepared for the knock-on effects of releasing all this [insecure] code into production … there are just not enough of us [security researchers] to handle this.”
@@ 7.2.2 For Platforms
Educate Users. Prior research in S&P has shown that even motivated users often struggle to make safe choices without clear, accessible guidance [110]. Building on these insights, we recommend that platforms develop educational resources to help users understand their operations and data-handling practices. These resources (e.g., tutorials, FAQ pages) can address common misconceptions, provide guidance on best practices for protecting personal data (e.g., adjusting privacy settings), and explain how user data is used for model training. By providing clear, targeted education, platforms can foster a more informed user base.
@@ 7.3 Limitations and Future Work
Given that our findings reflect specific concerns from professionals, developers, and other users using these platforms for specific applications, future work could explore users’ S&P concerns and attitudes in specific contexts and domains. By examining sector-specific issues, we can identify unique challenges and develop targeted recommendations for users and stakeholders in those fields. Similarly, future research should investigate the influence of geopolitical and cultural factors on user attitudes toward conversational AI platforms to understand regional differences and develop culturally sensitive solutions and policies.
@@ 8 Conclusion
Acknowledgments
We thank our anonymous reviewers and shepherd for providing us with valuable feedback. This work has been partially supported by UCI Academic Senate Council on Research, Computing and Libraries (CORCL) Award. Any findings, conclusions, and recommendations expressed in this paper are those of the authors only.
@@ Appendix B Keywords for Seed Corpus Creation
Table III shows the list of keywords for filtering posts.
@@ Appendix C S&P Classification Results
TABLE III: Keywords used to filter candidate posts for manual annotation. We applied regex-based matching to capture phrase variations (e.g., r’delete(?:my|your)?data’ to match “delete my data” and “delete your data”). Certain terms like access were matched as standalone words to avoid unrelated results (e.g., accessibility).
@@
Figure 4: Key events in the development and adoption of conversational AI platforms, including product and feature releases (green), policy and regulatory changes (yellow), corporate developments (navy), and privacy and security events (red).
Appendix D S&P-Related Posts Over Time
Figure 5 shows the weekly volume of S&P-related posts over time, averaging 344 posts per week.
Figure 5: Weekly counts of S&P-related posts.
@@ Appendix E Multi-Class Classification Results
Table IV shows the performance of the multi-class classification of S&P-related posts using GPT-4o [73].